### v1

# Mixture of Low Rank Adaptation with Partial Parameter Sharing for Time Series Forecasting

Abstract
Multi-task forecasting has become the standard approach for time-series forecasting (TSF). However, we show that it suffers from an Expressiveness Bottleneck, where predictions at different time steps share the same representation, leading to unavoidable errors even with optimal representations. To address this issue, we propose a two-stage framework: first, pre-train a foundation model for one-step-ahead prediction; then, adapt it using step-specific LoRA modules. This design enables the foundation model to handle any number of forecast steps while avoiding the expressiveness bottleneck. We further introduce the Mixture-of-LoRA (MoLA) model, which employs adaptively weighted LoRA experts to achieve partial parameter sharing across steps. This approach enhances both efficiency and forecasting performance by exploiting interdependencies between forecast steps. Experiments show that MoLA significantly improves model expressiveness and outperforms state-of-the-art time-series forecasting methods. Code is available at https://anonymous.4open.science/r/MoLA-BC92.
1 Introduction
Time-series forecasting (TSF) involves predicting future sequences based on observed ones and is widely applied in domains such as finance (e.g., stock price prediction (fintsb, )), meteorology (e.g., weather forecasting (weather_app, ; weather_app2, )), and manufacturing (e.g., process monitoring (energy_app, )). To construct effective TSF models, two questions warrant investigation: (1) How to extract informative representations from historical sequences; (2) How to generate accurate predictions using the extracted representations.
Recent research has predominantly focused on learning representations from historical sequences, where the key is developing neural network architectures that can effectively capture the temporal dynamics within time-series data. Exemplary works in this area include recurrent neural networks (deepar, ), convolutional neural networks (Timesnet, ; micn, ), and graph neural networks (FourierGNN, ). Current progress is characterized by an ongoing debate between Transformers and simple linear models. Transformers, equipped with self-attention mechanisms, provide superior scalability (itransformer, ; PatchTST, ; fredformer, ). In contrast, linear models, which encapsulate temporal dynamics using linear layers, are straightforward to implement and often demonstrate strong performance (DLinear, ; FBM, ; sparsetsf, ). These advancements showcase the rapid evolution of representation learning in the field of time-series data.
Despite significant progress in time-series representation learning, methods for generating predictions from these representations remain relatively underdeveloped. Specifically, current approaches predominantly adopt the multi-task forecasting (MT-F) paradigm, which employs a learnable linear layer to transform acquired representations into multi-step forecasts. In this work, we reveal that this approach is inherently constrained by an expressiveness bottleneck. Specifically, forecasts for different time steps are restricted to being linear combinations of shared representation bases. This limitation significantly constrains model capacity, leading to inevitable forecasting errors regardless of the quality of the learned representations.
To mitigate the expressiveness bottleneck, we propose a two-stage framework for constructing time-series forecasting models. It involves first training a single-step forecast model, and then adapting it to other forecast horizons by training step-specific LoRA (Low-Rank Adaptation) modules. However, directly applying this approach neglects the interdependencies between different forecast steps, resulting in huge computational costs and suboptimal performance. To mitigate these issues, we introduce the Mixture-of-LoRA (MoLA) model. This approach shares partial parameters among step-specific LoRA modules, which effectively reduces computational cost and enhances forecasting performance.
Our main contributions are summarized as follows:
  • We formalize the expressiveness bottleneck in the prevalent MT-F paradigm, which limits the capacity of forecasting models. We further propose a two-stage framework to construct time-series forecasting models free from this bottleneck.
  • We introduce MoLA to implement the two-stage framework. Building on the standard LoRA technique, MoLA enables partial parameter sharing among LoRA modules across different forecast steps, leveraging interdependencies between these steps to enhance both efficiency and accuracy.
  • We validate the efficacy of MoLA through extensive experiments, demonstrating its substantial superiority over state-of-the-art methods trained from scratch under the same forecasting lengths.
2 Preliminaries
In this work, we consider the time-series forecasting (TSF) problem: predicting future observations from historical data. Given a time-series dataset $X$ with $\mathrm{D}$ covariates, where $X_{n}$ denotes the observation at the $n$-th step, we define two key elements (box2015time, ): (1) Historical sequence $L=[X_{n-\mathrm{L}+1},\ldots,X_{n}]\in\mathbb{R}^{\mathrm{L}\times\mathrm{D}}$, where $\mathrm{L}$ is the historical window length; (2) Label sequence $Y=[X_{n+1},\ldots,X_{n+\mathrm{T}}]\in\mathbb{R}^{\mathrm{T}\times\mathrm{D}}$, where $\mathrm{T}$ is the forecast horizon. Based on these elements, TSF can be described as estimating $\mathbb{E}[Y|L]$—the expected label sequence conditioned on the past (nguyen2024scaling, ; ghimire2024two, ).
Most TSF models follow a two-stage architecture. The encoder $g_{\mathrm{e}}$ extracts informative representations from historical sequence: $R=g_{\mathrm{e}}(L)$; the decoder $g_{\mathrm{d}}$ then transforms $R$ to generate forecasts: $\hat{Y}=g_{\mathrm{d}}(R)$. While both stages are essential, recent research has mainly advanced encoder designs, developing diverse architectures such as convolutional neural networks (Timesnet, ; micn, ), graph neural networks (FourierGNN, ), and Transformers (itransformer, ; Transformer, ). In contrast, the decoder is understudied despite its importance.
There are two mainstream decoding approaches for TSF. Early methods adopt the autoregressive forecasting (AR-F) approach, where a linear layer with a single output produces one-step-ahead predictions, and the forecast sequence is generated sequentially by feeding previous outputs back as inputs. However, this approach suffers from error accumulation: errors in earlier steps interfere with later predictions. Modern TSF methods favor the multi-step forecasting (MT-F) paradigm (Informer, ; DLinear, ; itransformer, ). As illustrated in Figure 2(a), a linear layer with $\mathrm{T}$ outputs is applied to $R$ to generate the entire forecast sequence at once. This direct approach reduces error accumulation and eliminates the sequential computation overhead, making it the preferred choice in current TSF models.
3 Proposed Method
3.1 Motivation
MT-F has been a pervasive approach for generating forecasts from learned representations (itransformer, ; Timesnet, ; DLinear, ), as it mitigates error accumulation and eliminates the sequential computation overhead (Informer, ). Here, we formulate the MT-F workflow and identify an expressiveness bottleneck limiting its capacity.
Let $R\in\mathbb{R}^{\mathrm{L}\times\mathrm{D}}$ be the encoder output. MT-F generates the forecast sequence using a linear layer with learnable weights $W\in\mathbb{R}^{\mathrm{T}\times\mathrm{L}}$ and $b\in\mathbb{R}^{\mathrm{T}}$:
$\left[\hat{Y}_{1},\hat{Y}_{2},\ldots,\hat{Y}_{\mathrm{T}}\right]=\left[W_{1},W_{2},\ldots,W_{\mathrm{T}}\right]R+[b_{1},b_{2},...,b_{\mathrm{T}}].$ (1)
where $\hat{Y}_{t}=W_{t}R+b_{t}$ is the prediction for the $t$-th future step. This immediately assumes that different tasks (i.e., forecast steps) share the same optimal representation ($R$), such that applying different linear weights yields accurate forecasts. However, in practice, the optimal representation for each step may differ substantially, especially when the forecast horizon is large (e.g., $\mathrm{T}=720$). This mismatch results in an expressiveness bottleneck: regardless of encoder quality, using a shared representation across all forecast steps leads to an unavoidable modeling error. As shown in Theorem 3.1, this error is strictly positive when $\mathrm{L}+1<\mathrm{T}$ — a typical setting in the long-term TSF task.
Theorem 3.1 (Expressiveness Bottleneck).
Let $\bar{W}=[W\ b]\in\mathbb{R}^{\mathrm{T}\times(\mathrm{L}+1)}$ be the parameters in the MT-F’s linear decoder, $Y\in\mathbb{R}^{\mathrm{T}\times\mathrm{D}}$ be the label sequence; the minimum attainable estimation error is
$\|\epsilon\|_{F}^{2}=\sum_{t=\mathrm{rank}(\bar{W})+1}^{\mathrm{T}}\|U_{t}^{\top}Y\|_{2}^{2},$ (2)
where $\bar{W}=U\Sigma V^{\top}$ is the singular value decomposition of $\bar{W}$, $\mathrm{rank}(\bar{W})\leq\min\{\mathrm{T},\mathrm{L}+1\}$, and $\{U_{i}\}_{i=\mathrm{rank}+1}^{\mathrm{T}}$ form an orthonormal basis for the null space of $\bar{W}$. Notably, this error is independent of the representation $R$ provided by encoder.
(a) Synthetic data
(b) T=1
(c) T=16
(d) T=32
Figure 1: Visualization of representations generated with different forecasting step.
Case Study.
To illustrate the expressiveness bottleneck, we conduct a case study on a synthetic dataset. We set the historical window length to $\mathrm{L}=16$. The encoder is a two-layer perceptron with hidden sizes 16 and 2; the decoder is a linear layer with a single output. For each forecast step, we train a separate single-output model and visualize the encoder outputs in Figure 1 (b)-(d). The results show that the optimal representations differ significantly across steps, demonstrating that sharing a single representation across all forecast steps leads to unavoidable modeling errors.
Given the decoder’s critical role and the limitations of current approaches, there is a clear need for a decoding strategy that can overcome the expressiveness bottleneck. At the same time, the primary advantage of MT-F—parameter sharing across forecast steps—has proven beneficial for accuracy and should be preserved. To this end, two key questions warrant investigation: How to design a decoding strategy that avoids the expressiveness bottleneck while maintaining parameter sharing? Does addressing this bottleneck actually improve forecasting performance?
(a) MT-F
(b) MF
(c) MoLA
Figure 2: Visualization of MT-F, LoRA and MoLA approaches to generate multi-step forecasts. Gray blocks denote identical encoder components. Purple blocks represent decoding strategies. Rectangles with varying transparencies indicate different expert matrices in MoLA.
3.2 A Two-Stage Framework for Breaking Expressiveness Bottleneck
Motivated by Theorem 3.1, which shows that the error term disappears when the linear layer’s output dimension is 1, we advocate a two-stage framework to address the expressiveness bottleneck. We first pre-train a foundation model with a single output for one-step-ahead prediction and subsequently adapt it to multiple forecasting steps. The adaptation follows Low-Rank Adaptation (LoRA), which excels at adapting foundation models to different tasks without hampering pre-training performance.
Pre-training.
We begin by training a foundation model for one-step prediction. Given the historical sequence $L\in\mathbb{R}^{\mathrm{L}\times\mathrm{D}}$, the process can be described as follows:
$R=g_{\mathrm{e}}(L),\quad\hat{Y}_{1}=g_{\mathrm{d}}(R)=WR+b,\quad\mathcal{L}_{\mathrm{pre}}=\|Y_{1}-\hat{Y}_{1}\|_{2}^{2},$ (3)
where the encoder $g_{\mathrm{e}}$ extracts features $R$, and the one-step-ahead prediction $\hat{Y}_{1}\in\mathbb{R}^{\mathsf{D}}$ is generated using a single-output linear projection with weights $W\in\mathbb{R}^{1\times\mathrm{L}}$ and bias $b\in\mathbb{R}$. The pre-training objective is the MSE on the one-step prediction.
Adaptation.
After pre-training for one-step-ahead prediction ($\mathrm{T}=1$), we adapt the model for multi-step forecasting ($\mathrm{T}\geq 2$). To this end, we introduce LoRA modules into selected encoder linear layers, while freezing the weights of the original encoder and decoder. Specifically, for a selected linear layer with weight matrix $M\in\mathbb{R}^{\mathrm{d}_{\mathrm{out}}\times\mathrm{d}_{\mathrm{in}}}$, we modify the weights for each step $t$ as follows:
$M^{(t)}=M+B^{(t)}A^{(t)},$ (4)
where $B^{(t)}\in\mathbb{R}^{\mathrm{d}_{\mathrm{out}}\times r}$ and $A^{(t)}\in\mathbb{R}^{r\times\mathrm{d}_{\mathrm{in}}}$ are LoRA matrices, with rank $r<\min(\mathrm{d}_{\mathrm{out}},\mathrm{d}_{\mathrm{in}})$. For each forecast step $t$, the modified encoder $g_{\mathrm{e}}^{(t)}$ generates predictions as follows:
$R^{(t)}=g_{\mathrm{e}}^{(t)}(L),\quad\hat{Y}_{t}=g_{\mathrm{d}}(R^{(t)})=WR+b,\quad\mathcal{L}_{\mathrm{ada}}^{(t)}=\|Y_{t}-\hat{Y}_{t}\|_{2}^{2},$ (5)
where the adaptation objective is the MSE on the $t$-step prediction. Notably, only the LoRA matrices ($B^{(t)},A^{(t)}$) are optimized to minimize the adaptation objective, while the rest of the foundation model remains frozen. This process is repeated for each prediction step $1<t\leq\mathrm{T}$, adapting the foundation model to support diverse forecast steps.
Discussion.
The framework overcomes the expressiveness bottleneck by ensuring both the pretrained and adapted TSF models use a single output, thereby eliminating the error term as shown in Theorem 3.1. Moreover, during the adaptation stage, the pre-trained weights remain frozen, preventing any interference with the foundation model. This property has a critical implication for TSF: adapting to long-range forecasts does not compromise the performance on short-term predictions.
3.3 Mixture of Low-rank Adaptation
Time series data exhibits strong autocorrelation, inducing correlations across multiple forecasting steps. However, the above approach treats each forecast step independently by assigning separate LoRA matrices to generate representation and using a single-output layer for prediction. This leads to two limitations: (1) the number of adaptations increases linearly with the forecast horizon, resulting in high computational cost; and (2) neglecting inter-step correlations can impair forecasting accuracy.
In this section, we propose MoLA, a modified LoRA technique that leverages label correlations for improved performance and efficiency. According to Theorem 3.1, the expressiveness bottleneck vanishes as long as the number of outputs does not exceed $\mathrm{L+1}$; thus, single-step output is not strictly required. Building on this insight, we partition the $\mathrm{T}$-step label sequence into $\mathrm{K}$ contiguous segments, each containing $\mathrm{S}=\mathrm{T}/\mathrm{K}$ steps. The foundation model’s output dimension is then increased from 1 to $\mathrm{S}$, allowing each adaptation to adapt $\mathrm{S}$ future steps jointly, and thereby reducing the number of required adaptations. This segment-based adaptation exploits intra-segment label dependencies to improve accuracy and efficiency, while bypassing expressiveness bottleneck as long as $\mathrm{S}\leq\mathrm{L+1}$.
To further exploit label correlation across segments, we enable partial parameter sharing across different LoRA modules via a mixture-of-experts mechanism. Specifically, suppose we have $\mathrm{P}$ expert matrices, each denoted as $B^{(p)}$ and $A^{(p)}$. For a selected linear layer in the encoder with original weights $M$, the adaptation for segment $k$ is given by
$M^{(k)}=M+\sum_{p=1}^{\mathrm{P}}\Delta^{(p)}_{k}\cdot B^{(p)}A^{(p)},$ (6)
where $\Delta_{k}=[\Delta^{(1)}_{k},\Delta^{(2)}_{k},\ldots,\Delta^{\mathrm{(P)}}_{k}]$ is a learnable, normalized weight vector that adaptively combines the $\mathrm{P}$ expert matrices for each segment. This design enables expert matrices to be shared across segments, allowing the model to leverage inter-segment label correlations and enrich supervision signals for improved adaptation. A visual comparison among MT-F, LoRA, and MoLA is provided in Figure 2(a).
3.4 Overall Workflow
Algorithm 1 Mixture of Low-rank Adaptation for TSF.
In this section, we detail the procedure for applying MoLA to train TSF models, as outlined in Algorithm 1. The workflow begins by pre-training a foundation model for $\mathrm{S}$-step prediction (step 1). After pre-training, the model parameters are frozen (step 2). Next, we inject adaptation modules into the backbone: for each of the $\mathrm{K}$ segments, a segment-specific weight vector and $\mathrm{P}$ shared LoRA expert modules are introduced. For each segment, the mixing weights are learned to adaptively combine the shared LoRA experts, as defined in (6) (step 5). During fine-tuning, the segment-specific weight vector and segment-shared LoRA modules are optimized using the corresponding segment’s ground truth, while the backbone remains frozen (steps 6-8). And during inference, the segment-specific adapters are applied to the frozen foundation model to generate forecasts for each segment, which are then concatenated to produce the final estimations.
4 Experiments
To validate the effectiveness of our MoLA approach for TSF, we conduct a comprehensive empirical evaluation across six key dimensions:
  1. 1.
    Performance: How does MoLA perform compared to the current state-of-the-art? In Section 4.2, we benchmark MoLA against state-of-the-art baselines using public datasets to assess its overall performance.
  2. 2.
    Superiority: Is MoLA more effective than other forecasting paradigms? Section 4.3 provides a comparative analysis of MoLA versus AR-F and MT-F paradigms, demonstrating its advantages in forecasting accuracy.
  3. 3.
    Generality: Can MoLA be applied to other forecasting models? In Section 4.4, we assess the versatility of MoLA by applying it to various forecasting models, illustrating its broad applicability.
  4. 4.
    Flexibility: Does MoLA accommodate different fine-tuning modules besides LoRA? Section 4.4 explores MoLA’s flexibility by substituting LoRA with alternative fine-tuning modules, demonstrating its adaptability in implementation.
  5. 5.
    Sensitivity: How sensitive is MoLA to hyperparameter changes? In Section 4.5, we conduct a sensitivity analysis of the hyperparameter $\mathrm{K}$, showing that MoLA maintains robust performance across a wide range of parameter settings.
4.1 Experimental Setup
Datasets.
We utilize a diverse set of benchmark datasets for time-series forecasting, including ETT (with 4 subsets), ECL, Traffic, Weather, and PEMS (with 2 subsets), following the experimental settings used in (Autoformer, ) and (itransformer, ). Each dataset is split chronologically into training, validation, and test sets to preserve temporal order. Detailed descriptions of each dataset, including their characteristics and split ratios, can be found in Appendix D.1.
Baselines.
Our baselines consist of a wide range of established models from the time-series forecasting domain: Transformer (Transformer, ), Informer (Informer, ), Autoformer (Autoformer, ), FEDformer (fedformer, ), iTransformer (itransformer, ), DLinear (DLinear, ) and FreTS (FreTS, ), TimesNet (Timesnet, ) and TCN (tcn, ).
Implementation details.
To ensure consistency and fairness, we faithfully reproduced all baseline models using training scripts from the official TimesNet repository [^1] https://github.com/thuml/Time-Series-Library.. Optimization was performed using the Adam optimizer Adam , with learning rates selected from $\{0.0001,0.0005,0.001\}$. The forecasting horizons were set to 96, 192, 336, and 720 for the ETT, ECL, Traffic, and Weather datasets, and 12, 24, 36, and 48 for the PEMS datasets. We adopted mean square error (MSE) and mean absolute error (MAE) as primary evaluation metrics, in line with standard practices in the field.
To integrate the MoLA paradigm into existing models, we adhered to the original hyperparameter configurations of each baseline during the pre-training phase. Forecasting horizons of the pre-trained models were set to $\mathrm{T}/\mathrm{K}$, where $\mathrm{K}\in\{2,3,4,6\}$. In the fine-tuning phase, we only tuned the rank $r$ in the LoRA expert module, the learning rate $\eta$, and the number of LoRA experts, ensuring minimal deviation from the original model structures. All experiments were conducted on Intel(R) Xeon(R) Platinum 8383C CPUs and NVIDIA RTX 4090 GPUs. Additional implementation details are provided in Appendix D.2.
Table 1: Multi-step forecasting performance.
Models
MoLA
iTransformer
FreTS
TimesNet
TiDE
DLinear
FEDformer
Autoformer
Informer
Transformer
TCN
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
0.400
0.406
0.415
0.416
0.407
0.415
0.413
0.418
0.419
0.419
0.404
0.407
0.440
0.451
0.596
0.517
0.887
0.693
0.943
0.733
0.891
0.632
ETTm2
0.287
0.330
0.294
0.335
0.335
0.379
0.297
0.332
0.358
0.404
0.344
0.396
0.302
0.348
0.326
0.366
1.256
0.801
1.322
0.814
3.411
1.432
ETTh1
0.443
0.441
0.449
0.447
0.488
0.474
0.478
0.466
0.628
0.574
0.462
0.458
0.441
0.457
0.476
0.477
1.064
0.806
0.993
0.788
0.763
0.636
ETTh2
0.377
0.401
0.390
0.410
0.550
0.515
0.413
0.426
0.611
0.550
0.558
0.516
0.430
0.447
0.478
0.483
4.358
1.719
3.296
1.419
3.325
1.445
ECL
0.174
0.265
0.176
0.267
0.209
0.297
0.214
0.307
0.251
0.344
0.225
0.319
0.229
0.339
0.228
0.339
0.335
0.416
0.274
0.367
0.617
0.598
Traffic
0.425
0.284
0.428
0.286
0.552
0.348
0.535
0.309
0.760
0.473
0.673
0.419
0.611
0.379
0.637
0.399
0.727
0.404
0.680
0.376
1.001
0.652
Weather
0.266
0.286
0.281
0.302
0.255
0.299
0.262
0.288
0.271
0.320
0.265
0.317
0.311
0.361
0.349
0.391
0.595
0.541
0.632
0.552
0.584
0.572
PEMS03
0.112
0.222
0.116
0.226
0.146
0.257
0.118
0.223
0.316
0.370
0.233
0.344
0.174
0.302
0.501
0.513
0.137
0.241
0.126
0.233
0.666
0.634
PEMS08
0.138
0.236
0.159
0.258
0.174
0.277
0.154
0.245
0.319
0.378
0.294
0.377
0.232
0.322
0.630
0.572
0.319
0.314
0.249
0.266
0.713
0.629
  • Note: We fix the input length as 96 following the established benchmarks (itransformer, ; Timesnet, ). Bold typeface highlights the top performance for each metric, while underlined text denotes the second-best results. The results are averaged over prediction lengths (96, 192, 336 and 720), with full results in Table 6.
4.2 Overall Performance
The performance of our proposed modularized fine-tuning approach on the MSTF task is presented in Table 1. We use iTransformer as the base model $g_{\theta}$ and apply it to different forecasting horizons using the two-stage MoLA paradigm. Overall, MoLA significantly enhances the performance of iTransformer. For example, on the ETTm1 dataset, MoLA reduces the MSE of iTransformer by 0.015. Similar improvements are observed across other datasets, underscoring the effectiveness of the MoLA paradigm in modularized and personalized fine-tuning for different forecasting horizons.
Importantly, MoLA not only improves iTransformer’s performance but also enables it to surpass models that previously outperformed iTransformer on certain datasets and metrics, like PEMS08 with MSE and MAE. This suggests that the gains achieved by MoLA go beyond architectural designs alone, emphasizing the importance of modularized and personalized fine-tuning in addressing the expressiveness bottleneck often encountered in the MT-F paradigm.
Figure 3: Visualization of forecast sequence generated with and without MoLA under two snapshots.
Showcases.
To further illustrate the improvements provided by MoLA, we visualize forecast sequences for two snapshots from the ETTm2 dataset with a forecasting horizon of $\mathrm{T}=336$ in Figure 3. While the MT-F paradigm follows the general trends of the true label sequence, it struggles with capturing sharp peaks, leading to misaligned forecasts across different horizons. In contrast, MoLA mitigates this issue by employing horizon-specific fine-tuning, resulting in more accurate predictions that track both trends and sharp peaks across multiple horizons.
4.3 Comparative Studies
Table 2: Comparative study results.
ETTh1 Weather
Paradigm MSE $\Delta$ MAE $\Delta$ MSE $\Delta$ MAE $\Delta$
96
AR-F
0.454 - 0.441 - 0.235 - 0.264 -
MT-F
0.390 14.10% $\downarrow$ 0.410 7.03% $\downarrow$ 0.201 14.47% $\downarrow$ 0.247 6.44% $\downarrow$
MoLA
0.379 16.52% $\downarrow$ 0.400 9.30% $\downarrow$ 0.173 26.38% $\downarrow$ 0.211 20.08% $\downarrow$
192
AR-F
0.525 - 0.484 - 0.272 - 0.293 -
MT-F
0.443 15.62% $\downarrow$ 0.441 8.88% $\downarrow$ 0.250 8.09% $\downarrow$ 0.283 3.41% $\downarrow$
MoLA
0.436 16.95% $\downarrow$ 0.432 10.74% $\downarrow$ 0.246 9.56% $\downarrow$ 0.280 4.44% $\downarrow$
336
AR-F
0.580 - 0.518 - 0.368 - 0.357 -
MT-F
0.480 17.24% $\downarrow$ 0.457 11.78% $\downarrow$ 0.302 17.93% $\downarrow$ 0.317 11.20% $\downarrow$
MoLA
0.477 17.76% $\downarrow$ 0.456 11.97% $\downarrow$ 0.277 24.73% $\downarrow$ 0.296 17.09% $\downarrow$
720
AR-F
0.620 - 0.550 - 0.600 - 0.444 -
MT-F
0.484 21.94% $\downarrow$ 0.479 12.91% $\downarrow$ 0.370 38.33% $\downarrow$ 0.362 18.47% $\downarrow$
MoLA
0.480 22.58% $\downarrow$ 0.478 13.09% $\downarrow$ 0.367 38.83% $\downarrow$ 0.356 19.82% $\downarrow$
  • Note: $\Delta$ denotes the relative error improvement compared to iTransformer with AR-F paradigm.
This section presents a comparative analysis of the MoLA paradigm against the traditional AR-F and MT-F paradigms, using iTransformer as the base model. The evaluation covers two datasets (ETTh1 and Weather) and four different forecasting horizons, with the results summarized in Table 2.
Across all horizons, MoLA consistently outperforms both AR-F and MT-F paradigms. While MT-F reduces error accumulation, which is a known issue with AR-F, MoLA goes a step further by mitigating the expressiveness bottleneck of MT-F. By introducing modularized and personalized fine-tuning for each forecasting horizon, MoLA enhances the forecast accuracy, particularly for longer forecasting horizons where performance gains are more pronounced. These findings demonstrate that MoLA is more than just an incremental improvement over MT-F, it represents a meaningful advancement in fine-tuning techniques for TSF to overcome the expressiveness bottleneck.
4.4 Generalization Studies
In this section, we explore the generality of the MoLA paradigm to enhance varying forecasting models and encompass existing fine-tuning techniques.
Generalization to Fine-tuning Techniques.
Table 3: Generalization results.
Variants MT-F MoLA-Ada MoLA-IA3 MoLA-R MoLA
Metrics MSE $\Delta$ MAE $\Delta$ MSE $\Delta$ MAE $\Delta$ MSE $\Delta$ MAE $\Delta$ MSE $\Delta$ MAE $\Delta$ MSE $\Delta$ MAE $\Delta$
ETTh2 96 0.301 - 0.349 - 0.292 2.93$\downarrow$ 0.342 2.10$\downarrow$ 0.304 0.88$\uparrow$ 0.346 0.84$\downarrow$ 0.293 2.53$\downarrow$ 0.342 2.13$\downarrow$ 0.293 2.57$\downarrow$ 0.341 2.25$\downarrow$
192 0.382 - 0.402 - 0.384 0.55$\uparrow$ 0.398 1.08$\downarrow$ 0.381 0.17$\downarrow$ 0.397 1.34$\downarrow$ 0.380 0.54$\downarrow$ 0.398 1.00$\downarrow$ 0.377 1.25$\downarrow$ 0.397 1.32$\downarrow$
336 0.430 - 0.434 - 0.420 2.31$\downarrow$ 0.431 0.78$\downarrow$ 0.418 2.88$\downarrow$ 0.428 1.48$\downarrow$ 0.422 1.81$\downarrow$ 0.433 0.23$\downarrow$ 0.421 2.06$\downarrow$ 0.429 1.22$\downarrow$
720 0.447 - 0.455 - 0.430 3.87$\downarrow$ 0.446 1.97$\downarrow$ 0.437 2.23$\downarrow$ 0.447 1.68$\downarrow$ 0.428 4.25$\downarrow$ 0.444 2.44$\downarrow$ 0.418 6.49$\downarrow$ 0.438 3.67$\downarrow$
Avg 0.390 - 0.410 - 0.382 2.05% $\downarrow$ 0.402 1.95% $\downarrow$ 0.385 1.28% $\downarrow$ 0.405 1.22% $\downarrow$ 0.381 2.31% $\downarrow$ 0.402 1.98% $\downarrow$ 0.377 3.33% $\downarrow$ 0.401 2.20% $\downarrow$
Weather 96 0.201 - 0.247 - 0.202 0.43$\uparrow$ 0.246 0.39$\downarrow$ 0.204 1.69$\uparrow$ 0.249 0.80$\uparrow$ 0.202 0.40$\uparrow$ 0.245 1.00$\downarrow$ 0.173 13.78$\downarrow$ 0.211 14.71$\downarrow$
192 0.250 - 0.283 - 0.248 0.83$\downarrow$ 0.279 1.23$\downarrow$ 0.250 0.14$\downarrow$ 0.281 0.63$\downarrow$ 0.248 0.90$\downarrow$ 0.279 1.45$\downarrow$ 0.246 1.43$\downarrow$ 0.280 1.02$\downarrow$
336 0.302 - 0.317 - 0.299 0.89$\downarrow$ 0.317 0.03$\downarrow$ 0.300 0.60$\downarrow$ 0.315 0.66$\downarrow$ 0.280 7.28$\downarrow$ 0.297 6.44$\downarrow$ 0.277 8.28$\downarrow$ 0.296 6.63$\downarrow$
720 0.370 - 0.362 - 0.361 2.41$\downarrow$ 0.351 3.04$\downarrow$ 0.365 1.40$\downarrow$ 0.356 1.66$\downarrow$ 0.357 3.57$\downarrow$ 0.349 3.63$\downarrow$ 0.367 0.86$\downarrow$ 0.356 1.60$\downarrow$
Avg 0.281 - 0.302 - 0.278 1.07% $\downarrow$ 0.299 0.99% $\downarrow$ 0.280 0.36% $\downarrow$ 0.300 0.66% $\downarrow$ 0.272 3.20% $\downarrow$ 0.293 2.98% $\downarrow$ 0.266 5.34% $\downarrow$ 0.286 5.30% $\downarrow$
  • Note: $\Delta$ denotes the relative error improvement compared to iTransformer with MT-F paradigm.
We implement MoLA by replacing the LoRA modules with two alternative fine-tuning techniques: Adapter (Adapter, ) (MoLA-Ada) and IA3 (IA3, ) (MoLA-IA3), to show its support of existing fine-tuning techniques. Both Adapter and IA3 are well-established parameter-efficient fine-tuning technologies. We also introduce a variant that applies step-specific modularized fine-tuning with standard LoRA modules (MoLA-R) for comparison. Detailed illustrations of their technical differences and parameter settings are provided in Appendix E.3. The results in Table 3 indicate that the devised variants exhibit comparable improvements over the MT-F paradigm, affirming MoLA’s flexibility in integrating diverse fine-tuning strategies. The standard MoLA, enabling parameter sharing across different tasks with the MoE mechanism, outperforms the variants since it leverages the fruitful inter-step information while maintaining model efficiency.
Generalization to Forecasting Models.
We incorporate MoLA into several well-established TSF models: iTransformer, Autoformer, Informer, and Transformer. The results, averaged across different prediction lengths (96, 192, 336, 720) and accompanied by 95% confidence intervals, are presented in Figure 4. Overall, MoLA improves the performance of these models over the traditional MT-F paradigm. And notably, Autoformer and Informer benefit significantly from the incorporation of MoLA, showing a relative MSE improvement of over 4% on both the ECL and Weather datasets. These results demonstrate the generality and broad applicability of MoLA, reinforcing its potential as a plug-and-play strategy for enhancing various neural forecasting models in TSF.
(a) ECL with MSE.
(b) ECL with MAE
(c) Weather with MSE.
(d) Weather with MAE.
Figure 4: Benefit of incorporating MoLA in varying models, shown with colored bars for means over forecasting lengths (96, 192, 336, 720) and error bars for 95% confidence intervals.
4.5 Sensitivity Studies
Sensitivity on Segment Number.
Table 4: Results of varying segment number.
ETTh1 Weather
$\mathrm{K}$ MSE $\Delta$ MAE $\Delta$ MSE $\Delta$ MAE $\Delta$
MT-F 0.390 - 0.410 - 0.201 - 0.247 -
1 0.390 0.09% $\downarrow$ 0.410 0.21% $\downarrow$ 0.201 0.02% $\downarrow$ 0.246 0.69% $\downarrow$
2 0.383 2.02% $\downarrow$ 0.403 1.79% $\downarrow$ 0.202 0.25% $\uparrow$ 0.246 0.61% $\downarrow$
3 0.389 0.33% $\downarrow$ 0.409 0.24% $\downarrow$ 0.200 0.54% $\downarrow$ 0.244 1.23% $\downarrow$
4 0.379 2.93% $\downarrow$ 0.400 2.47% $\downarrow$ 0.201 0.03% $\downarrow$ 0.245 1.07% $\downarrow$
8 0.391 0.14% $\uparrow$ 0.408 0.70% $\downarrow$ 0.173 13.99% $\downarrow$ 0.211 14.68% $\downarrow$
16 0.390 0.10% $\downarrow$ 0.407 0.79% $\downarrow$ 0.176 12.57% $\downarrow$ 0.214 13.53% $\downarrow$
32 0.388 0.66% $\downarrow$ 0.403 1.81% $\downarrow$ 0.174 13.43% $\downarrow$ 0.213 14.02% $\downarrow$
48 0.389 0.27% $\downarrow$ 0.405 1.37% $\downarrow$ 0.198 1.50% $\downarrow$ 0.244 1.47% $\downarrow$
96 0.389 0.31% $\downarrow$ 0.404 1.61% $\downarrow$ 0.181 9.92% $\downarrow$ 0.221 10.84% $\downarrow$
  • Note: $\Delta$ denotes the relative error improvement compared to iTransformer with MT-F paradigm. The results are generated with forecasting length fixed at 96.
In this section, we investigate the impact of the segment number $\mathrm{K}$ on the performance of MoLA. We conduct experiments on the ETTh1 and Weather datasets by varying $\mathrm{K}$ while keeping $\mathrm{T}=96$, and the results are presented in Table 4. First, MoLA outperforms MT-F across nearly all values of $\mathrm{K}$. Secondly, smaller segment sizes (e.g., $\mathrm{K}=4$ for ETTh1 and $\mathrm{K}=8$ for the Weather dataset) achieve the best performance, demonstrating the effectiveness of the segmentation strategy in enhancing MoLA’s forecasting accuracy since it enables the utilization of correlations in neighboring steps. Therefore, although ideally each prediction length would have its own set of LoRA weights, i.e., $\mathrm{K}=\mathrm{T}$, the segmentation strategy offers a flexible solution that reduces computational complexity while simultaneously improving forecasting performance.
Sensitivity on Other Hyperparameters.
We then examine the sensitivity of MoLA paradigm to three key hyperparameters: the rank $r$ of the LoRA expert modules, the learning rate $\eta$ used during fine-tuning and the number of LoRA expert modules. Experiments on ETTh1 and Weather datasets (prediction lengths 192 and 336) reveal distinct patterns, as shown in Figure 5. First, for rank $r$ (tested over $\{4,8,16,32,64\}$), both excessively small and large ranks degrade performance, with the optimal value balancing flexibility and overfitting risks. For instance, on ETTh1, $r=8$ achieves minimal MSE values of 0.441 (192) and 0.484 (336), while higher ranks overfit and lower ranks lack expressiveness. Second, learning rate $\eta$ exhibits sharper sensitivity: increasing $\eta$ from 0.0002 to 0.001 on ETTh1 reduces MSE from 0.459 to 0.440 for 192-step forecasts, aligning with prior observations on its critical role in convergence (LoRA-sensi, ). Finally, the number of LoRA experts shows stable performance unless it severely mismatches the segment count $\mathrm{K}$. For example, on Weather (192-step), expanding experts from 2 to 10 only reduces MSE by 0.001, suggesting robustness to moderate variations. These trends highlight the necessity of calibrating $r$ and $\eta$ precisely, while the expert count offers flexibility within practical bounds. Extended results across forecasting lengths are provided in Appendix E.4.
(a) Varying $r$ on ETTh1.
(b) Varying $\eta$ on ETTh1.
(c) Varying $\mathrm{P}$ on ETTh1.
(d) Varying $r$ on Weather.
(e) Varying $\eta$ on Weather.
(f) Varying $\mathrm{P}$ on Weather.
Figure 5: Performance given varying rank $r$, learning rate $\eta$ and the number of experts $\mathrm{P}$.
5 Conclusion
In this work, we address the expressiveness bottleneck in time-series forecasting, a limitation that hinders accurate multi-step prediction with existing multi-task paradigm. We identify that conventional linear decoders, by sharing representations across different forecast steps, are inherently inadequate for capturing step-specific dynamics, resulting in unavoidable forecast errors. To handle this challenge, we propose a two-stage framework: first, pre-training a foundation model for one-step-ahead prediction, and then adapting it for multi-step forecasting using step-specific Low-Rank Adaptation (LoRA) modules. Building on this design, we further introduce the Mixture-of-LoRA (MoLA) model, which leverages segment-based adaptation and adaptively weighted LoRA experts to achieve partial parameter sharing across forecast steps. This approach enables MoLA to flexibly balance expressiveness and efficiency, while effectively exploiting interdependencies between forecast steps. Extensive experiments across diverse datasets and backbone models demonstrate that MoLA consistently improves forecasting accuracy and outperforms state-of-the-art time-series forecasting methods, highlighting the promise of modularized fine-tuning and partial parameter sharing for multi-horizon forecasting.
Limitations and Future Work. This study primarily focuses on adaptation using LoRA expert modules for fine-tuning the $\mathrm{S}$-step foundation model. However, the parameter cost of LoRA still scales with model size and forecast horizon. Exploring alternative parameter-efficient adaptation techniques, such as prompt-tuning or adapter-tuning, may further enhance the scalability and applicability of the proposed framework. Additionally, in this work, the segmentation of forecast steps is manually specified. Automatically determining segment boundaries based on data-driven analysis of time-series patterns and periodicity represents a promising direction for further optimizing the MoLA framework.
Appendix A Related Work
A.1 Time Series Forecasting Modeling
Time-series forecasting (TSF) modeling generally involves encoding historical sequences and decoding temporal representations. To exploit temporal dynamics in the encoding phase, various deep learning backbones have been developed, generally grouped into four main categories: RNN-based (e.g., SegRNN (segrnn, )), CNN-based (e.g., TimesNet (Timesnet, )), GNN-based (e.g., MAGNN (magnn, )), MLP-based, and Transformer-based methods. Recent debates focus on MLP versus Transformer architectures, where MLPs (e.g., DLinear (DLinear, ), TSMixer (TSMixer2, )) are efficient but limited in handling complex temporal patterns, whereas Transformers (e.g., PatchTST (PatchTST, ), iTransformer (itransformer, )) excel in temporal encoding but are computationally intensive. To better capture intricate temporal patterns, specialized designs such as series decomposition (e.g., Autoformer (Autoformer, )) and multiperiodicity analysis (e.g., FiLM (film, )) have been proposed, addressing seasonal and mixed period forecasting, respectively. Recent innovations explore frequency-domain representations for temporal patterns, exemplified by FedFormer (fedformer, ) which employs frequency-domain attention score computation to enhance both efficiency and effectiveness. This paradigm demonstrates remarkable adaptability across architectures, including Transformers (fedformer, ; Autoformer, ), MLPs (FreTS, ), and GNNs (FourierGNN, ; stemgnn, ), establishing frequency-domain analysis as a versatile plug-and-play component for temporal modeling.
Despite significant advancements in the encoding phase, the decoding phase remains underexplored in TSF modeling. The dominant multi-task forecasting (MT-F) paradigm employs simplistic affine transformations to map temporal representations to multi-step predictions. While computationally efficient, this approach inherently suffers from an expressiveness bottleneck, where predictions are confined to linear combinations of shared representation bases. Our theoretical analysis reveals that such constrained decoding mechanisms induce systematic error accumulation and suboptimal prediction expressiveness, even with well-encoded temporal patterns, exposing fundamental limitations in conventional MT-F paradigm.
A.2 Modularized Fine-tuning
Fine-tuning has emerged as a pivotal technique for adapting pre-trained models to downstream tasks, initially gaining prominence in natural language processing and computer vision (CV-FT, ). Modern approaches have evolved into modular frameworks categorized into three paradigms: Adapter-based, Selection-based, and LoRA-based methods. The Adapter-based paradigm introduces lightweight task-specific modules between pre-trained layers, preserving original parameters while enabling domain adaptation (serial-adapter, ; parallel_adapter, ; CoDA, ; mera, ). In contrast, Selection-based methods employ parameter masking strategies to identify critical subnetworks for task-specific tuning (PaFi, ; SPT, ; FAR, ; cross_attn, ). The LoRA-based paradigm marks a significant technical evolution in parameter-efficient fine-tuning by introducing Low-Rank Adaptation (LoRA) (LoRA, ; note_on_lora, ), which augments certain layers of a pre-trained model with trainable low-rank matrices. Instead of updating the full set of model parameters, LoRA only optimizes a small, low-rank decomposition of the weight update, thereby drastically reducing memory and computational overhead. This design allows for efficient adaptation to new tasks with minimal storage and enables fast switching between tasks by maintaining separate LoRA weights. LoRA’s effectiveness and versatility have been further demonstrated through updated matrix decomposition (loretta, ; adalora, ), quantization (qlora, ; loftq, ), and ranking adaptation (sora, ; cora, ), which reflects a broader trend toward modularized and scalable LoRA applications.
In the context of TSF, LoRA-based fine-tuning has demonstrated promising capabilities from two main perspectives. On one hand, (TSF-FT5, ) introduces channel-aware LoRA, leveraging low-rank adaptation to capture channel dependencies. On the other hand, a series of studies (TSF-FT1, ; TSF-FT2, ; TSF-FT3, ; TSF-FT4, ) investigate the impact of LoRA within time series foundation models, focusing on how LoRA facilitates efficient adaptation and task transfer in time series forecasting. Though these explorations are promising, they largely overlook two critical aspects: the expressiveness bottleneck we have identified in TSF, and the untapped potential of LoRA for overcoming this limitation. Our work diverges by specifically addressing the MT-F paradigm’s expressiveness bottleneck through innovative integration of temporal segmentation and Mixture-of-LoRA enhanced decoding, ensuring both model efficiency and expressiveness.
Appendix B Complexity Analysis
In this section, we conduct a parameter count analysis and evaluate the running cost of MoLA through empirical investigation. Take iTransformer (itransformer, ) as the base model, let the number of layers in iTransformer be $\mathrm{N}_{l}$, the hidden dimension of the attention layer be $d_{\text{m}}$, and the hidden dimension of the FFN layer be $d_{\text{ff}}$. Then, the number of parameters introduced by MoLA is given by:
$\mathrm{N}_{\text{MoLA}}$ $=\mathrm{N}_{l}\times 2\times(\mathrm{N}_{\text{LoRA}}\times\mathrm{P}+\mathrm{N}_{\text{Weight}}\times\mathrm{K})$ (7)
$=\mathrm{N}_{l}\times 2\times((d_{\text{m}}\times r+r\times d_{\text{ff}})\times\mathrm{P}+\mathrm{P}\times\mathrm{K}),$
where $\mathrm{N}_{\text{LoRA}}$ and $\mathrm{N}_{\text{Weight}}$ are the number of parameters introduced by LoRA expert module and learnable weights for each expert, respectively.
Consider that iTransformer is a standard Transformer architecture with dimension permutation in the input sequence, its number of parameters can be given by:
$\mathrm{N}_{\text{iTrans}}$ $=\mathrm{N}_{l}\times(\mathrm{N}_{\text{Att}}+\mathrm{N}_{\text{FFN}}+\mathrm{N}_{\text{LN}})$ (8)
$=\mathrm{N}_{l}\times(4\times d_{\text{m}}^{2}+2\times d_{\text{m}}\times d_{\text{ff}}+4\times d_{\text{m}}),$
where $\mathrm{N}_{\text{Att}}$, $\mathrm{N}_{\text{FFN}}$ and $\mathrm{N}_{\text{LN}}$ are the number of parameters introduced by self attention layer, feedforward layer and layer normalization layer, respectively.
The ratio between the two parameter counts is therefore:
$\frac{\mathrm{N}_{\text{MoLA}}}{\mathrm{N}_{\text{iTrans}}}=\frac{\mathrm{P}\times(d_{\text{m}}\times r+r\times d_{\text{ff}}+\mathrm{K})}{2\times d_{\text{m}}\times(d_{\text{m}}+d_{\text{ff}}+1)}.$ (9)
For example, with $\mathrm{P}=4$, $\mathrm{K}=6$, $d_{\text{m}}=512$, $d_{\text{ff}}=1024$, and $r=8$, the ratio $\frac{\mathrm{N}_{\text{MoLA}}}{\mathrm{N}_{\text{iTrans}}}$ is approximately 0.047, indicating a relevant small increase in model size. Figure 6 shows MoLA’s empirical running costs with varying $r$, $\mathrm{P}$, and $\mathrm{K}$, capturing the sum time for weight matrix adaptation and base layer forward pass. Our results confirm that the additional computational duration imposed by MoLA is lower than 1ms, with a small increase as the hyperparameters grow. Therefore, MoLA does not compromise the model’s efficiency while effectively improving the model’s performance.
(a) Varying $r$ on ETTh1.
(b) Varying $\mathrm{P}$ on ETTh1.
(c) Varying $\mathrm{K}$ on ETTh1.
(d) Varying $r$ on Weather.
(e) Varying $\mathrm{P}$ on Weather.
(f) Varying $\mathrm{K}$ on Weather.
Figure 6: Running time of MoLA given varying rank $r$, the number of experts $\mathrm{P}$ and the number of segments $\mathrm{K}$.
Appendix C Theoretical Justification
Theorem C.1 (Expressiveness Bottleneck).
Let $\bar{W}=[W\ b]\in\mathbb{R}^{\mathrm{T}\times(\mathrm{L}+1)}$ be the parameters in the MT-F’s linear decoder, $Y\in\mathbb{R}^{\mathrm{T}\times\mathrm{D}}$ be the label sequence; the minimum attainable estimation error is
$\|\epsilon\|_{F}^{2}=\sum_{t=\mathrm{rank}(\bar{W})+1}^{\mathrm{T}}\|U_{t}^{\top}Y\|_{2}^{2},$ (10)
where $\bar{W}=U\Sigma V^{\top}$ is the singular value decomposition of $\bar{W}$, $\mathrm{rank}(\bar{W})\leq\min\{\mathrm{T},\mathrm{L}+1\}$, and $\{U_{i}\}_{i=\mathrm{rank}+1}^{\mathrm{T}}$ form an orthonormal basis for the null space of $\bar{W}$. Notably, this error is independent of the representation $R$ provided by encoder.
Proof.
Consider the least squares estimation, the aim of linear decoder is to find $\bar{R}=[R^{\top},1]^{\top}\in\mathbb{R}^{(\mathrm{L}+1)\times 1}$ such that:
$\hat{\bar{R}}=\arg\min_{\bar{R}}\|Y-\bar{W}\bar{R}\|_{2}^{2}.$
To solve this optimization problem, we set the gradient of the cost function with respect to $\bar{R}$ to zero:
$\frac{\partial}{\partial\bar{R}}\|Y-\bar{W}\bar{R}\|_{2}^{2}=-2\bar{W}^{\top}(Y-\bar{W}\bar{R})=0.$
Assuming $\bar{W}^{\top}\bar{W}$ is invertible, we have:
$\bar{R}=(\bar{W}^{\top}\bar{W})^{-1}\bar{W}^{\top}Y.$
Then, the estimation error is:
$\epsilon=Y-\hat{Y}=\left(I-\bar{W}(\bar{W}^{\top}\bar{W})^{-1}\bar{W}^{\top}\right)Y=(I-P)Y,$
where $P=\bar{W}(\bar{W}^{\top}\bar{W})^{-1}\bar{W}^{\top}$ is known as the projection matrix onto the column space of $\bar{W}$. Therefore, $I-P$ projects onto the orthogonal complement of the column space of $\bar{W}$.
Let the Singular Value Decomposition (SVD) of $\bar{W}$ be
$\bar{W}=U\Sigma V^{\top},$
where $U\in\mathbb{R}^{\mathrm{T}\times\mathrm{T}}$ is orthogonal, $\Sigma\in\mathbb{R}^{\mathrm{T}\times(\mathrm{L}+1)}$ is diagonal, and $V\in\mathbb{R}^{(\mathrm{L}+1)\times(\mathrm{L}+1)}$ is orthogonal.
Partition $\Sigma$ and $U$ as
$\Sigma=\begin{bmatrix}\Sigma_{o}&0\\ 0&0\end{bmatrix},\quad U=[U_{o}\quad U_{n}],$
where $\Sigma_{o}\in\mathbb{R}^{o\times o}$ contains the positive singular values, $U_{o}\in\mathbb{R}^{\mathrm{T}\times o}$ contains the corresponding left singular vectors, and $U_{n}\in\mathbb{R}^{\mathrm{T}\times(\mathrm{T}-o)}$ contains the left singular vectors corresponding to zero singular values, and $o=\mathrm{rank}(\bar{W})$.
The projection matrix onto the column space of $\bar{W}$ is
$P=\bar{W}(\bar{W}^{\top}\bar{W})^{-1}\bar{W}^{\top}=U_{o}U_{o}^{\top}.$
Therefore, the estimation error is
$\epsilon=(I-P)Y=U_{n}U_{n}^{\top}Y.$
The minimum attainable estimation error is
$\|\epsilon\|_{F}^{2}=\|U_{n}^{\top}Y\|_{2}^{2}=\sum_{t=\mathrm{rank}(\bar{W})+1}^{\mathrm{T}}\left\|U_{t}^{\top}Y\right\|_{2}^{2}.$
Theorem C.2 (Variance Reduction of MoLA).
$\mathcal{L}_{\text{MoLA}}$ has a smaller variance than $\mathcal{L}_{\text{MT-F}}$,
$\operatorname{Var}(\mathcal{L}_{\text{MoLA}})\leq\operatorname{Var}(\mathcal{L}_{\text{MT-F}}).$ (11)
Proof.
Consider the total loss expressed as the average of losses over the prediction horizons:
$\mathcal{L}=\frac{1}{\mathrm{T}}\sum_{t=1}^{\mathrm{T}}\mathcal{L}_{t},$
where $\mathcal{L}_{t}$ is the MSE loss for the $t$-th prediction horizon.
The variance of the total loss is:
[…]
Under modularized fine-tuning, the covariances between different $\mathcal{L}_{t}$ decrease because the LoRA modules introduce horizon-specific parameters, reducing parameter sharing. Let $\Delta\operatorname{Cov}(\mathcal{L}_{t},\mathcal{L}_{s})=\operatorname{Cov}_{\text{MT-F}}(\mathcal{L}_{t},\mathcal{L}_{s})-\operatorname{Cov}_{\text{MoLA}}(\mathcal{L}_{t},\mathcal{L}_{s})\geq 0$. The variance difference is then:
$\operatorname{Var}(\mathcal{L}_{\text{MT-F}})-\operatorname{Var}(\mathcal{L}_{\text{MoLA}})=\frac{2}{\mathrm{T}^{2}}\sum_{1\leq t<s\leq\mathrm{T}}\Delta\operatorname{Cov}(\mathcal{L}_{t},\mathcal{L}_{s})\geq 0.$
Therefore,
$\operatorname{Var}(\mathcal{L}_{\text{MoLA}})&=\operatorname{Var}(\mathcal{L}_{\text{MT-F}})-\frac{2}{\mathrm{T}^{2}}\sum_{1\leq t<s\leq\mathrm{T}}\Delta\operatorname{Cov}(\mathcal{L}_{t},\mathcal{L}_{s})\\ &\leq\operatorname{Var}(\mathcal{L}_{\text{MT-F}}).\end{aligned}$
The equation realize only when all the LoRA modules do not contribute to the improvement of the prediction performance with corresponding time step. However, we didn’t recognize that phenomenon in our experiments.
Appendix D Reproduction Details
D.1 Dataset Descriptions
Table 5: Detailed dataset descriptions. D denotes the number of variates. Forecast Length denotes the prediction lengths investigated in this dataset. Frequency denotes the sampling interval of time points. Train, Validation, Test denotes the number of samples employed in each split. The taxonomy and statistic are aligned with the recent works (Timesnet, ; itransformer, ).
Dataset D Forecast Length Split Ratio Frequency Domain
ETTm1 7 96,192,336,720 34465/11521/11521 15mins Electricity
ETTm2 7 96,192,336,720 34465/11521/11521 15mins Electricity
ETTh1 7 96,192,336,720 8545/2881/2881 Hourly Electricity
ETTh2 7 96,192,336,720 8545/2881/2881 Hourly Electricity
ECL 321 96,192,336,720 18317/2633/5261 Hourly Electricity
Traffic 862 96,192,336,720 12185/1757/3509 Hourly Transportation
Weather 21 96,192,336,720 36792/5271/10540 10mins Weather
PEMS03 358 12,24,36,48 15617/5135/5135 5mins Transportation
PEMS08 170 12,24,36,48 10690/3548/265 5mins Transportation
The datasets used in this study span a variety of domains and time resolutions, each with distinct characteristics that are well-suited for evaluating time-series forecasting:
  • ETT (Informer, ): This dataset consists of data from 7 key variables related to electricity transformers, collected between July 2016 and July 2018. It includes four subsets: ETTh1 and ETTh2, which are recorded hourly, and ETTm1 and ETTm2, which are recorded every 15 minutes.
  • ECL (Electricity Consumption Load) (Autoformer, ): This dataset contains hourly electricity consumption data from 321 clients, offering insights into power usage patterns over time.
  • Traffic (Autoformer, ): This dataset comprises hourly road occupancy rates collected by 862 sensors deployed across the freeways in the San Francisco Bay Area. The data spans the period from January 2015 to December 2016, reflecting traffic conditions over time.
  • Weather (Autoformer, ): This dataset includes 21 meteorological variables, recorded every 10 minutes throughout 2020 at the Max Planck Biogeochemistry Institute’s weather station. It provides a comprehensive set of climate-related factors for forecasting purposes.
  • PEMS (SCINet, ): This dataset consists of public traffic data from the California highway network, with recordings collected every 5 minutes. We utilize two subsets in this study: PEMS03 and PEMS08, which are frequently adopted in traffic forecasting benchmarks.
For all datasets, the data processing and division into training, validation, and test sets follow the protocols established by TimesNet (Timesnet, ) and iTransformer (itransformer, ), ensuring a consistent chronological split to avoid data leakage. The standardized lookback window is set at $96$ for the ETT, ECL, Traffic, Weather, and PEMS datasets. Prediction horizons vary across datasets, with forecasting lengths of $96$, $192$, $336$, and $720$ for the first five datasets, and shorter horizons of $12$, $24$, $36$, and $48$ for the PEMS subsets. Detailed specifications of each dataset can be found in Table 5.
D.2 Implementation Details
The baseline models in this study were carefully reproduced using training scripts from the TimesNet Repository (Timesnet, ), ensuring full reproducibility verification. All models were trained using the Adam optimizer (Adam, ), with learning rates selected from $\{10^{-3},5\times 10^{-4},10^{-4}\}$ based on minimizing the MSE loss. A batch size of 32 was maintained consistently across all experiments. Training was performed for a maximum of 10 epochs, with an early stopping criterion triggered if no improvement in validation performance was observed for 3 consecutive epochs.
For the experiments integrating MoLA into existing forecasting models, we strictly adhered to the original hyperparameter settings as outlined in the respective publications during the pre-training phase. The forecasting horizon of pre-trained models were set to $\mathrm{T}/\mathrm{K}$, where $\mathrm{K}\in\{2,3,4,6\}$. Pre-training was limited to a maximum of 5 epochs, with early stopping applied after 2 epochs without improvement. During the fine-tuning phase, only the rank $r$ in the LoRA expert module, the learning rate $\eta$ and the number of LoRA expert modules $\mathrm{P}$ were tuned, as these parameters are crucial for adjusting the differences in weight magnitudes between the base models and the LoRA expert modules. Fine-tuning was performed to minimize the MSE averaged over all prediction lengths, with forecasting segments selected from range $\{2,3,4,6\}$. The current results sufficiently demonstrate the effectiveness of MoLA, showing that its efficacy is not dependent on highly specific hyperparameter configurations.
Appendix E More Experimental Results
E.1 Overall Performance
Table 6: Full results on the multi-step forecasting task. The length of history window is set to 96 for all baselines. Avg indicates the results averaged over forecasting lengths: T=96, 192, 336 and 720.
Models
MoLA
iTransformer
FreTS
TimesNet
TiDE
DLinear
FEDformer
Autoformer
Informer
Transformer
TCN
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
ETTm1
96
0.330
0.364
0.346
0.379
0.339
0.374
0.338
0.379
0.364
0.387
0.345
0.372
0.389
0.427
0.468
0.463
0.633
0.560
0.591
0.549
0.887
0.613
192
0.378
0.392
0.392
0.400
0.382
0.397
0.389
0.400
0.398
0.404
0.381
0.390
0.402
0.431
0.573
0.509
0.736
0.625
0.704
0.629
0.877
0.626
336
0.414
0.416
0.427
0.422
0.421
0.426
0.429
0.428
0.428
0.425
0.414
0.414
0.438
0.451
0.596
0.527
1.061
0.787
1.171
0.861
0.890
0.636
720
0.480
0.453
0.494
0.461
0.485
0.462
0.495
0.464
0.487
0.461
0.473
0.451
0.529
0.498
0.749
0.569
1.119
0.801
1.307
0.893
0.911
0.653
Avg
0.400
0.406
0.415
0.416
0.407
0.415
0.413
0.418
0.419
0.419
0.404
0.407
0.440
0.451
0.596
0.517
0.887
0.693
0.943
0.733
0.891
0.632
ETTm2
96
0.181
0.262
0.184
0.266
0.190
0.282
0.185
0.264
0.207
0.305
0.195
0.294
0.194
0.284
0.240
0.319
0.541
0.581
0.317
0.408
3.125
1.345
192
0.247
0.307
0.257
0.315
0.260
0.329
0.254
0.307
0.290
0.364
0.283
0.359
0.264
0.324
0.300
0.349
0.527
0.558
1.069
0.758
3.130
1.350
336
0.312
0.347
0.315
0.351
0.373
0.405
0.314
0.345
0.377
0.422
0.384
0.427
0.319
0.359
0.339
0.375
1.126
0.797
1.325
0.869
3.185
1.375
720
0.407
0.403
0.419
0.409
0.517
0.499
0.434
0.413
0.558
0.524
0.516
0.502
0.430
0.424
0.423
0.421
2.828
1.268
2.576
1.223
4.203
1.658
Avg
0.287
0.330
0.294
0.335
0.335
0.379
0.297
0.332
0.358
0.404
0.344
0.396
0.302
0.348
0.326
0.366
1.256
0.801
1.322
0.814
3.411
1.432
ETTh1
96
0.379
0.400
0.390
0.410
0.399
0.412
0.422
0.433
0.479
0.464
0.396
0.410
0.377
0.418
0.423
0.441
0.920
0.745
0.796
0.691
0.767
0.633
192
0.436
0.432
0.443
0.441
0.453
0.443
0.465
0.457
0.521
0.503
0.449
0.444
0.421
0.445
0.498
0.485
0.998
0.781
0.813
0.699
0.739
0.619
336
0.477
0.456
0.480
0.457
0.503
0.475
0.492
0.470
0.659
0.603
0.487
0.465
0.468
0.472
0.506
0.496
1.091
0.812
1.181
0.876
0.717
0.613
720
0.480
0.478
0.484
0.479
0.596
0.565
0.532
0.502
0.893
0.736
0.516
0.513
0.500
0.493
0.477
0.487
1.247
0.887
1.182
0.885
0.828
0.678
Avg
0.443
0.441
0.449
0.447
0.488
0.474
0.478
0.466
0.628
0.574
0.462
0.458
0.441
0.457
0.476
0.477
1.064
0.806
0.993
0.788
0.763
0.636
ETTh2
96
0.293
0.341
0.301
0.349
0.350
0.403
0.320
0.364
0.400
0.440
0.343
0.396
0.347
0.391
0.383
0.424
2.340
1.220
2.072
1.140
3.171
1.364
192
0.377
0.397
0.382
0.402
0.472
0.475
0.409
0.417
0.528
0.509
0.473
0.474
0.430
0.443
0.557
0.511
6.284
2.078
5.081
1.814
3.222
1.398
336
0.421
0.429
0.430
0.434
0.564
0.528
0.449
0.451
0.643
0.571
0.603
0.546
0.469
0.475
0.470
0.481
4.824
1.853
3.564
1.475
3.306
1.452
720
0.418
0.438
0.447
0.455
0.815
0.654
0.473
0.474
0.874
0.679
0.812
0.650
0.473
0.480
0.501
0.515
3.985
1.724
2.469
1.247
3.599
1.565
Avg
0.377
0.401
0.390
0.410
0.550
0.515
0.413
0.426
0.611
0.550
0.558
0.516
0.430
0.447
0.478
0.483
4.358
1.719
3.296
1.419
3.325
1.445
ECL
96
0.147
0.237
0.148
0.239
0.189
0.277
0.171
0.273
0.237
0.329
0.210
0.302
0.200
0.315
0.199
0.315
0.315
0.398
0.252
0.352
0.688
0.621
192
0.163
0.254
0.167
0.258
0.193
0.282
0.188
0.289
0.236
0.330
0.210
0.305
0.207
0.322
0.215
0.327
0.327
0.411
0.266
0.364
0.587
0.582
336
0.179
0.270
0.179
0.272
0.207
0.296
0.208
0.304
0.249
0.344
0.223
0.319
0.226
0.340
0.232
0.343
0.354
0.434
0.292
0.383
0.590
0.588
720
0.208
0.297
0.209
0.298
0.245
0.332
0.289
0.363
0.284
0.373
0.258
0.350
0.282
0.379
0.268
0.371
0.343
0.423
0.287
0.371
0.602
0.601
Avg
0.174
0.265
0.176
0.267
0.209
0.297
0.214
0.307
0.251
0.344
0.225
0.319
0.229
0.339
0.228
0.339
0.335
0.416
0.274
0.367
0.617
0.598
Traffic
96
0.396
0.271
0.397
0.272
0.528
0.341
0.504
0.298
0.805
0.493
0.697
0.429
0.577
0.362
0.609
0.385
0.698
0.390
0.686
0.385
1.451
0.744
192
0.416
0.278
0.418
0.279
0.531
0.338
0.526
0.305
0.756
0.474
0.647
0.407
0.603
0.372
0.633
0.400
0.697
0.386
0.679
0.377
0.842
0.622
336
0.425
0.283
0.432
0.286
0.551
0.345
0.540
0.310
0.762
0.477
0.653
0.410
0.615
0.378
0.637
0.398
0.715
0.397
0.663
0.361
0.844
0.620
720
0.464
0.304
0.467
0.305
0.598
0.367
0.570
0.324
0.719
0.449
0.694
0.429
0.649
0.403
0.668
0.415
0.797
0.443
0.693
0.381
0.867
0.624
Avg
0.425
0.284
0.428
0.286
0.552
0.348
0.535
0.309
0.760
0.473
0.673
0.419
0.611
0.379
0.637
0.399
0.727
0.404
0.680
0.376
1.001
0.652
Weather
96
0.173
0.211
0.201
0.247
0.184
0.239
0.178
0.226
0.202
0.261
0.197
0.259
0.221
0.304
0.284
0.355
0.383
0.438
0.332
0.383
0.610
0.568
192
0.246
0.280
0.250
0.283
0.223
0.275
0.227
0.266
0.242
0.298
0.236
0.294
0.275
0.345
0.313
0.371
0.415
0.449
0.634
0.539
0.541
0.552
336
0.277
0.296
0.302
0.317
0.272
0.316
0.283
0.305
0.287
0.335
0.282
0.332
0.338
0.379
0.359
0.393
0.618
0.551
0.656
0.579
0.565
0.569
720
0.367
0.356
0.370
0.362
0.340
0.363
0.359
0.355
0.351
0.386
0.347
0.384
0.408
0.418
0.440
0.446
0.963
0.726
0.908
0.706
0.622
0.601
Avg
0.266
0.286
0.281
0.302
0.255
0.299
0.262
0.288
0.271
0.320
0.265
0.317
0.311
0.361
0.349
0.391
0.595
0.541
0.632
0.552
0.584
0.572
PEMS03
12
0.067
0.173
0.069
0.175
0.083
0.194
0.082
0.188
0.117
0.225
0.122
0.245
0.123
0.248
0.239
0.365
0.122
0.226
0.107
0.209
0.632
0.606
24
0.095
0.205
0.098
0.210
0.127
0.241
0.110
0.216
0.233
0.320
0.202
0.320
0.160
0.287
0.492
0.506
0.129
0.233
0.121
0.227
0.655
0.626
36
0.127
0.240
0.131
0.243
0.169
0.281
0.133
0.236
0.380
0.422
0.275
0.382
0.191
0.321
0.399
0.459
0.143
0.249
0.133
0.243
0.678
0.644
48
0.159
0.270
0.164
0.275
0.204
0.311
0.146
0.251
0.536
0.511
0.335
0.429
0.223
0.350
0.875
0.723
0.153
0.255
0.144
0.253
0.699
0.659
Avg
0.112
0.222
0.116
0.226
0.146
0.257
0.118
0.223
0.316
0.370
0.233
0.344
0.174
0.302
0.501
0.513
0.137
0.241
0.126
0.233
0.666
0.634
PEMS08
12
0.079
0.181
0.085
0.189
0.095
0.204
0.110
0.209
0.121
0.231
0.152
0.274
0.175
0.275
0.446
0.483
0.268
0.281
0.213
0.236
0.680
0.607
24
0.114
0.218
0.131
0.236
0.150
0.259
0.142
0.239
0.232
0.326
0.245
0.350
0.211
0.305
0.488
0.509
0.296
0.302
0.238
0.256
0.701
0.622
36
0.158
0.256
0.182
0.282
0.202
0.305
0.167
0.258
0.379
0.428
0.344
0.417
0.250
0.338
0.532
0.513
0.340
0.327
0.263
0.277
0.727
0.637
48
0.203
0.290
0.236
0.323
0.250
0.341
0.195
0.274
0.543
0.527
0.437
0.469
0.293
0.371
1.052
0.781
0.373
0.345
0.283
0.295
0.746
0.648
Avg
0.138
0.236
0.159
0.258
0.174
0.277
0.154
0.245
0.319
0.378
0.294
0.377
0.232
0.322
0.630
0.572
0.319
0.314
0.249
0.266
0.713
0.629
$1^{\text{st}}$ Count
33
36
0
0
4
0
2
6
0
0
1
3
4
0
1
0
0
0
0
0
0
0
We present a comprehensive comparison of the multi-step forecasting task in Table 6. The iTransformer model serves as the base model for implementing the MoLA paradigm. Despite iTransformer’s initial performance gap compared to other state-of-the-art baseline models, integrating MoLA significantly enhances its forecasting accuracy.
Specifically, MoLA achieves the lowest MSE in 33 out of 45 cases and the lowest MAE in 36 out of 45 cases, illustrating its effectiveness in enhancing model performance. The improvements brought by MoLA are particularly noticeable on challenging datasets such as ETTm1 and Traffic, where capturing long-term dependencies is crucial. These results underscore the robustness and adaptability of the MoLA paradigm. While there are a few instances where MoLA does not achieve the top performance, this can be attributed to the inherent advantages of specific models in particular contexts. For example, FreTS shows competitive results on the Weather dataset, where its architectural design may better suit certain meteorological patterns. Nonetheless, the overall performance of MoLA demonstrates its strength as a fine-tuning framework for time-series forecasting, consistently mitigating the expressiveness bottleneck and delivering superior results across a wide range of forecasting tasks.
E.2 Comparative Studies
Table 7: Results on the multi-step forecasting task with AR-F paradigm. iTransformer is served as the base and single-step model. Avg indicates the results averaged over forecasting lengths: T=96, 192, 336 and 720 for ETT, ECL, Traffic and Weather dataset, T=12,24,36 and 48 for PEMS datasets.
Datasets
ETTm1
ETTm2
ETTh1
ETTh2
ECL
Traffic
Weather
PEMS03
PEMS08
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
AR-F
96
0.371
0.385
0.200
0.278
0.454
0.441
0.339
0.368
0.307
0.360
0.590
0.364
0.192
0.234
12
0.089
0.204
0.088
0.195
192
0.446
0.425
0.262
0.315
0.525
0.484
0.418
0.413
0.396
0.424
0.603
0.385
0.240
0.273
24
0.177
0.287
0.262
0.337
336
0.493
0.461
0.463
0.443
0.580
0.518
0.456
0.449
0.550
0.513
0.861
0.495
0.368
0.357
36
0.315
0.382
0.562
0.486
720
0.629
0.520
0.430
0.418
0.620
0.550
0.448
0.458
0.978
0.706
1.341
0.588
0.600
0.444
48
0.515
0.487
1.133
0.674
Avg
0.485
0.448
0.339
0.363
0.545
0.498
0.415
0.422
0.558
0.501
0.849
0.458
0.350
0.327
Avg
0.274
0.340
0.511
0.423
We present a detailed comparison of the multi-step forecasting task using the AR-F paradigm in Table 7, where iTransformer serves as both the base and single-step model. Overall, the AR-F paradigm demonstrates a significant issue with error accumulation as the forecasting horizon increases. This effect is particularly noticeable in datasets like PEMS08, where, for example, the MSE dramatically increases as the forecasting length extends from 12 to 48 time steps.
This behavior underscores the limitations of the AR-F paradigm for long-term forecasting tasks. As the model is required to recursively predict future values based on previous predictions, the compounding effect of errors leads to substantially degraded performance over longer horizons. This observation highlights the need for more robust solutions, such as our MoLA paradigm, which mitigates the error accumulation problem in AR-F and expressiveness bottleneck in MT-F by adopting a more modularized and horizon-specific fine-tuning approach, allowing for more stable and accurate long-term forecasts.
E.3 Generalization Studies
Figure 7: Visualization of common parameter-efficient fine-tuning strategies.
We visualize the three prominent PEFT methods: Adapter, LoRA and IA3 in Figure 7, showing the architectural modifications introduced by each PEFT technique within a typical transformer block. Adapter, shown in the upper left, introduces additional fully connected (FC) layers and a short-cut connection after the FFN layer. The Adapter module typically consists of a down-projection FC layer, followed by a non-linearity (often ReLU), an up-projection FC layer, and a residual connection. This approach provides a compact, trainable module that can adjust the model’s behavior for specific tasks without modifying the entire network. LoRA, illustrated on the lower left, modifies the FFN layer by adding low-rank matrices to the frozen weight matrix. LoRA decomposes the weight update into two low-rank matrices, enabling the model to learn task-specific adaptations in a parameter-efficient manner. IA3, shown on the right, works by applying learnable scaling factors to the key and value projections in the self-attention mechanism and to the hidden representations in the feed-forward network, allowing for fine-grained control over the model’s behavior with minimal additional parameters.
For a fair comparison, we matched the low-rank configuration of the Adapter and the LoRA modules with that of the LoRA expert in MoLA, applying both to the feed-forward network (FFN) layers. IA3, on the other hand, dynamically adjusts the weights of intermediate hidden vectors in both the FFN and attention layers.
E.4 Hyperparameter Sensitivity
In this section, we further analyze the sensitivity of MoLA to three key hyperparameters under different segment number settings: the low-rank value $r$ in the LoRA modules, the learning rate $\eta$, and the number of LoRA expert modules $\mathrm{P}$. The experiments are conducted using iTransformer across two datasets, ETTh1 and Weather, with the results visualized in Figure 8 and Figure 9.
The results generally indicate that both extremely low and high values of rank $r$ and learning rate $\eta$ lead to performance degradation. This pattern suggests that overly high ranks may lead to overfitting, while excessively low ranks may not provide sufficient flexibility for effective model adaptation. Similarly, very high learning rates can cause instability in training, while very low learning rates may result in slow convergence or getting stuck in suboptimal solutions. Interestingly, for the number of expert modules $\mathrm{P}$, MoLA exhibits high stability as long as $\mathrm{P}$ is not significantly lower than the number of segments $\mathrm{K}$. This robustness indicates that the model can effectively leverage multiple experts to capture diverse patterns across different forecasting segments.
These findings indicate that fine-tuning the hyperparameters for each specific forecasting horizon can further improve performance. However, our results demonstrate that even without exhaustive tuning, the MoLA paradigm delivers robust performance improvements, highlighting its flexibility and effectiveness in time-series forecasting tasks.
(a) Varying $r$ on ETTh1.
(b) Varying $\eta$ on ETTh1.
(c) Varying $\mathrm{P}$ on ETTh1.
Figure 8: Performance of iTransformer enhanced by MoLA given different low rank of LoRA modules $r$, learning rate $\eta$ and the number of experts $\mathrm{P}$. Different columns correspond to different number of forecasting length $\mathrm{T}$ (from left to right: 96, 192, 336, 720). The results are averaged on four forecasting segment number (2, 3, 4, 6) with shaded areas being 50% confidence intervals.
(a) Varying $r$ on Weather.
(b) Varying $\eta$ on Weather.
(c) Varying $\mathrm{P}$ on Weather.
Figure 9: Performance of iTransformer enhanced by MoLA given different low rank of LoRA modules $r$, learning rate $\eta$ and the number of experts $\mathrm{P}$. Different columns correspond to different number of forecasting length $\mathrm{T}$ (from left to right: 96, 192, 336, 720). The results are averaged on four forecasting segment number (2, 3, 4, 6) with shaded areas being 50% confidence intervals.
E.5 Extra Discussions
Table 8: Full results on the multi-step forecasting task with jointly fine-tuning weight and bias under MF paradigm. The length of history window is set to 96 for all baselines. Avg indicates the results averaged over forecasting lengths: T=96, 192, 336 and 720 for ETT, ECL, Traffic and Weather dataset, T=12,24,36 and 48 for PEMS datasets.
Datasets
ETTm1
ETTm2
ETTh1
ETTh2
ECL
Traffic
Weather
PEMS03
PEMS08
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MoLA-WB
96
0.337
0.368
0.182
0.263
0.378
0.400
0.297
0.345
0.147
0.237
0.394
0.270
0.201
0.245
12
0.070
0.175
0.079
0.181
192
0.382
0.394
0.246
0.306
0.434
0.431
0.379
0.398
0.163
0.254
0.420
0.282
0.246
0.279
24
0.095
0.205
0.114
0.218
336
0.419
0.420
0.315
0.350
0.477
0.456
0.420
0.431
0.182
0.273
0.640
0.426
0.280
0.297
36
0.127
0.239
0.158
0.256
720
0.483
0.456
0.415
0.407
0.495
0.488
0.432
0.448
0.211
0.300
0.585
0.352
0.367
0.356
48
0.158
0.269
0.202
0.290
Avg
0.405
0.409
0.289
0.332
0.446
0.444
0.382
0.405
0.176
0.266
0.549
0.353
0.274
0.294
Avg
0.112
0.222
0.138
0.236
Discussion on fine-tuning bias.
In this section, we extend our analysis by evaluating the impact of jointly fine-tuning both the weight and bias within the LoRA modules under the MoLA paradigm. The results, detailed in Table 8, cover a variety of datasets.
Overall, the findings indicate that the joint fine-tuning of weights and biases (MoLA-WB) results in performance degradation compared to fine-tuning weights alone under the MoLA paradigm. Specifically, MoLA-WB tends to increase the risk of overfitting, particularly in datasets with complex temporal dependencies, where the added non-low-rank parameters introduce excessive flexibility, making the model more difficult to optimize effectively. This is especially evident in the larger forecasting horizons, where the performance gap becomes more pronounced.
However, MoLA-WB generally improves the performance of the MT-F paradigm, suggesting that while the joint fine-tuning strategy introduces challenges for the MoLA framework, it may still offer some benefit in simpler or less modular fine-tuning strategies. The overall results affirm that focusing on weight fine-tuning alone is a more effective approach for leveraging the full potential of MoLA, ensuring better generalization and predictive accuracy across diverse forecasting tasks.
Discussion on Layer-wise Fine-tuning.
In this section, we explore the impact of layer-wise fine-tuning in the multi-layer iTransformer model across various datasets. Specifically, for the ETT dataset, we adaptd the first (MoLA-1) and second (MoLA-2) layers, while for the ECL, Traffic, Weather, and PEMS datasets, we extended the study to include the third layer (MoLA-3). The results are detailed in Tables 9 and Table 10.
From the experiments, we observe that layer-wise fine-tuning generally yields positive performance improvements over the MT-F paradigm in most cases, particularly for datasets like ETTm1, ETTm2, Weather, and ECL. In these datasets, fine-tuning a single layer was sufficient to capture the temporal dependencies effectively, leading to reduced MSE and MAE. For instance, in the ETT datasets, both MoLA-1 and MoLA-2 show competitive results compared to full-layer fine-tuning, indicating that focusing on specific layers can provide significant computational savings without sacrificing accuracy.
However, for more complex datasets such as Traffic and PEMS03, fine-tuning a single layer did not achieve results better than those obtained with the MT-F paradigm. This could be due to the disruption of key inter-layer interactions that are crucial for the hierarchical processing in Transformer-based models. These interactions are particularly important in datasets with complex temporal patterns or multiple variates, where adjustments in a single layer may not provide enough capacity to adapt to the nuances of the time-series data.
The results suggest that while layer-wise fine-tuning can be beneficial in reducing the computational overhead and maintaining high performance, it is dataset-dependent. In datasets with more complex structures, a more comprehensive fine-tuning strategy involving multiple layers or full-layer fine-tuning may be necessary to avoid underfitting and fully capture the temporal dependencies in the data.
Table 9: Full results on the multi-step forecasting task with layer-wise fine-tuning under MoLA paradigm.
Datasets
ETTm1
ETTm2
ETTh1
ETTh2
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MoLA-1
96
0.342
0.373
0.183
0.263
0.379
0.400
0.293
0.342
192
0.386
0.396
0.251
0.310
0.436
0.432
0.380
0.397
336
0.426
0.424
0.315
0.352
0.477
0.456
0.428
0.434
720
0.495
0.462
0.412
0.406
0.486
0.481
0.428
0.445
Avg
0.412
0.414
0.290
0.333
0.444
0.442
0.382
0.404
MoLA-2
96
0.335
0.369
0.181
0.262
0.377
0.398
0.296
0.343
192
0.382
0.395
0.246
0.306
0.434
0.431
0.384
0.402
336
0.414
0.418
0.308
0.344
0.480
0.459
0.432
0.438
720
0.480
0.454
0.411
0.405
0.498
0.490
0.430
0.445
Avg
0.403
0.409
0.286
0.329
0.447
0.445
0.385
0.407
Table 10: Full results on the multi-step forecasting task with layer-wise fine-tuning under MF paradigm.
Datasets
ECL
Traffic
Weather
PEMS03
PEMS08
Metrics
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MSE
MAE
MoLA-1
96
0.149
0.239
0.401
0.275
0.203
0.245
12
0.069
0.175
0.081
0.185
192
0.164
0.255
0.425
0.286
0.250
0.281
24
0.099
0.210
0.123
0.228
336
0.182
0.274
0.432
0.286
0.280
0.298
36
0.136
0.249
0.173
0.271
720
0.215
0.301
0.480
0.320
0.368
0.358
48
0.172
0.284
0.222
0.309
Avg
0.177
0.267
0.435
0.292
0.275
0.296
Avg
0.119
0.229
0.150
0.249
MoLA-2
96
0.149
0.240
0.401
0.276
0.203
0.246
12
0.069
0.175
0.082
0.185
192
0.163
0.254
0.423
0.285
0.248
0.280
24
0.099
0.209
0.122
0.228
336
0.182
0.274
0.431
0.286
0.281
0.298
36
0.135
0.248
0.174
0.272
720
0.212
0.299
0.479
0.319
0.369
0.358
48
0.171
0.281
0.218
0.305
Avg
0.176
0.267
0.433
0.292
0.275
0.295
Avg
0.118
0.228
0.149
0.247
MoLA-3
96
0.150
0.240
0.402
0.276
0.201
0.245
12
0.069
0.176
0.083
0.186
192
0.162
0.254
0.424
0.287
0.249
0.280
24
0.099
0.210
0.125
0.230
336
0.180
0.272
0.431
0.286
0.280
0.297
36
0.136
0.249
0.175
0.272
720
0.210
0.297
0.481
0.319
0.369
0.358
48
0.171
0.282
0.219
0.305
Avg
0.176
0.266
0.435
0.292
0.275
0.295
Avg
0.119
0.229
0.150
0.248
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect our paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Our limitations are discussed in conclusions.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All proofs of theoretical results are involved in Appendix.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We detail our training and evaluation protocols in the experimental setting section.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We use open access data, and the code is provided in an anonymous link.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We detail our training and evaluation protocols in the experimental setting section.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: In Figure 4 which are crucial for evaluating the efficacy of MoLA, we report the error bars to make the results more rigorous and comprehensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the type of compute workers, memory in Appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: Since it is an algorithm-oriented research, there is no societal impact of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not involve these issues.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The creators or original owners of assets (e.g., code, data, models), used in the paper, are properly credited. The license and terms of use are explicitly mentioned and properly respected.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not introduce new datasets and benchmarks in the paper.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We use open-access datasets and do not involve human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We use open-access datasets and do not involve human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We do not involve LLMs as any important, original, or non-standard components in this work.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 0.248
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: The main claims made in the abstract and introduction accurately reflect our paper’s contributions and scope.
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: Our limitations are discussed in conclusions.
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: All proofs of theoretical results are involved in Appendix.
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: We detail our training and evaluation protocols in the experimental setting section.
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: We use open access data, and the code is provided in an anonymous link.
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: We detail our training and evaluation protocols in the experimental setting section.
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: In Figure 4 which are crucial for evaluating the efficacy of MoLA, we report the error bars to make the results more rigorous and comprehensive.
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: We report the type of compute workers, memory in Appendix.
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: The research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics.
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: Since it is an algorithm-oriented research, there is no societal impact of the work performed.
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: This paper does not involve these issues.
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59.
-Justification: The creators or original owners of assets (e.g., code, data, models), used in the paper, are properly credited.
-The license and terms of use are explicitly mentioned and properly respected.
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: We do not introduce new datasets and benchmarks in the paper.
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: We use open-access datasets and do not involve human subjects.
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: We use open-access datasets and do not involve human subjects.
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: We do not involve LLMs as any important, original, or non-standard components in this work.
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.